# How to verify that a given process is a Lévy-Driven Ornstein-Uhlenbeck Process

### v1

@@ Abstract
11footnotetext: Department of Mathematics and Statistics, Rhodes College, Memphis, TN 38112, USA. Emails: abdelrazeqi@rhodes.edu, smirh-25@rhodes.edu, zhadi-25@rhodes.edu
@@ 1 Introduction
Between 2014 and 2018, Abdelrazeq et al. developed theories and results indicating how to test if a process is a Lévy-driven CAR(1) process [[1], [2], [3]]. Using these results, we develop a step-by-step methodological approach for how these findings can be used to verify the process and test its driving process distribution. The performance of these test statistics under the null hypothesis is illustrated via simulation studies. To demonstrate the power of the test, we computed the power under one alternative case. We then illustrate how to apply the methodology step-by-step to many economic and financial data examples.
@@ 3 Methodology
In this section, we outline the step-by-step methodology for testing different background processes. The approach and results are clearly presented to allow users to apply these techniques to their own models and data. See Sections 1 and 2 of the supplemental material for further explanation the methodology along with corresponding R code.
@@ 3.1 Step 1: Choosing the right estimator for $a$
We are concerned with the consistency of the test statistic $W_{\Delta_{1}\widehat{L}^{(M)}}(1)$ defined if we replace the parameter $a$ by an estimator. Thus, we must first choose the most accurate estimator for $a$.
@@
We need to consider an estimator for $a$ for the general second order driving process $L$ (i.e. $\mu\neq 0$). Consider the following estimator of $a$:
$\widehat{{a}}_{N}^{(M)}=\frac{\sum_{n=1}^{NM}\left(Y_{\frac{n-1}{M}}-Y_{\frac{n}{M}}\right)\left(Y_{\frac{n-1}{M}}-\overline{Y}\right)}{\frac{1}{M}\sum_{n=1}^{NM}\left(Y_{\frac{n-1}{M}}-\overline{Y}\right)^{2}}~{}~{}\text{where}~{}~{}\overline{Y}=\frac{1}{NM}\sum_{n=1}^{NM}Y_{\frac{n}{M}}.$ (3.1)
This estimator is based on the least squares approach as in [1]; hence we will refer to it as the least squares based estimator or LSB. This estimator is recommended for use when the driving process is an unspecified Lévy process, or when Brownian motion is expected to be involved in the driving process. Moreover, one should use this estimator in order to test if the driving process is Brownian motion or not. The explanation and R code for this estimator can be found in Section 1.4.1 of the supplemental material.
@@
This estimator is based on the highly efficient Davis-McCormick estimator as in [7]; hence we refer to it as Davis-McCormick based estimator or DMB estimator. This estimator is incredibly accurate as compared to the LSB estimator; however, it should not be used if Brownian motion is expected to be involved in the driving process. We have tested and recommend this estimator under Gamma, Inverse Gaussian, Beta, and any mixed combinations of such Lévy processes.
In general, the DMB estimator does not take negative values for $Y$. Hence, if $Y$ is negative at any point, then use the LSB estimator. The explanation and R code for this estimator can be found in Section 1.4.2 of the supplemental material.
Below in Tables 1-3, we observe the comparison of the accuracy of our two estimators in cases when the driving process is not Brownian motion, and we fix $\sigma=1$, $\mu=1$, $\eta=1$, $N=100$, $M=100$, and $\lambda=1$ in relevant cases.
@@
For the Gamma driven process, we observe that the DMB estimator is more accurate in estimation for all values of $a$.
@@
For the Inverse Gaussian driven process, we observe that the DMB estimator is more accurate in estimation for all values of $a$.
@@
For the mixed Inverse Gaussian and Gamma driven process, we observe that the DMB estimator is more accurate in estimation for all values of $a$.
We observe that the DMB estimator is more precise than the LSB estimator in most cases, yet both are relatively accurate in estimation for each background process. Moreover, the LSB estimator should be used when the driving process is Brownian motion or any unspecified driving process.
@@ 3.2 Step 2: Choosing N and M
Before calculating the recovered increments, one should choose $N$ and $M$ such that $N/M$ is small, c.f. [2]. $N$ can represent a day, a month, or a year where $M$ is the number of observations in each period accordingly. Also, we have observed that for small $a$ and $N$, the test has had some issues with the p-value being close to the nominal level. Hence, we recommend that N be more than 50.
@@ 3.3 Step 3: Recovering increments with estimated $a$
In order to proceed, we define the recovered increments using an estimator of $a$:
@@
See Section 1.7 of our supplemental material for R code alongside further explanation of the recovered increments with estimated $a$.
@@ 3.4 Step 4: Calculating the test statistic
We calculate $W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)$ as defined in Equation (2.8), then for the significance level $\alpha=0.05$, we reject the null hypothesis of uncorrelated increments when $|W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)|>1.96$. In this case, we reject the null hypothesis that the data is Lévy-driven for every background process. If $|W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)|<1.96$, then we fail to reject the null, and it is possible that Lévy-driven CAR(1) is a good model, so afterwards, one may want to check the background process. In this case, we recommend that you continue to Step 5 (see subsection 3.5). Refer to Section 1.8 of the supplemental material for how to calculate the test statistic in R.
Note that a good indication that a test is performing well under the null is that for a significance level of 0.05, and a number of simulations, 400 ($R=400$ in the following Tables), the p-value should be around $0.05\pm 0.02$, which is true in our most cases. It is acceptable to be off one out of twenty, by chance, by the binomial distribution.
@@
We have defined a Lévy-driven CAR(1) process, a method for recovering the observed increments with an estimated $a$, and the test statistic $W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)$. We now observe the performance of our verification process for various background processes (i.e. Brownian Motion, Gamma, Inverse Gaussian, Beta, and mixed Lévy processes). The tables for each background process display the p-value, see Equation (2.9), for different $N$, $M$, and estimated $a$ values. See Section 2.1 of the supplemental material for the performance of the various background processes along with corresponding R code.
@@
For the Brownian motion driven process, we observe p-values very near $\alpha=0.05$ for most combination of $N$, $M$, and $a$ values, and it seems it seems that the ratio $N/M$ has no effect. For $a=0.3$, and $a=0.5$, we observe the p-values are close to the nominal level, and we refer to [2] for further information on the effect of the LSB estimator on the result.
@@
For the Gamma driven process, we observe p-values very near $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it seems that the ratio $N/M$ has no effect.
@@
For the Inverse Gaussian driven process, we observe p-values very near $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it seems that the ratio $N/M$ has no effect.
@@
For the mixed Inverse Gaussian and Gamma driven process, we observe p-values very near $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it seems that the ratio $N/M$ has no effect.
Our results emphasizes the strength of the verification process, as we fail to reject the null hypothesis of uncorrelated increments for many Lévy-driven background processes and with varying combinations of $N$, $M$, and $a$ values. See Section 2 of the supplemental material for testing the performance of various background processes.
@@ 3.5 Step 5: Testing for the driving process
If the null hypothesis that the driving process is Lévy is not rejected, then we want to test whether the driving process is Brownian motion or not. Another interesting test is whether the driving process is some other specific process, say Gamma, Inverse Gaussian, or any other unspecified process. Here we recommend the following two procedures, generally taken from [3] and [10], and modified to our verification process, in order to accomplish these objectives.
@@
If we want to test if the driving process is BM or not, then we follow these steps.
@@
    Recover the estimated increments, as in Equation (3.3), using the LSB estimator.
@@
    Find a one bootstrap sample from the calculated estimated increments
@@
    Find the sample mean and standard deviation of the bootstrapped sample found in step 2
@@
    Conduct KS test with the sample mean and standard deviation calculated in step 3.
This procedure should only be used if the process that you are testing for is Brownian motion, see [6].
@@
Taken from [10], if we want to test whether the driving process is some process other than Brownian motion, i.e., Gamma, Inverse Gaussian, or any other unspecified process, we follow these steps:
@@
    Recover the estimated increments, as in Equation 3.3, using the appropriate estimator (LSB or DMB)
@@
    Assuming the recovered increments follow a particular distribution $F(.;\theta)$, estimate the parameters for this distribution from the recovered increments
@@
    Calculate the test statistic $D_{N}$ for the $N$ recovered increments $Z_{i}$, where $i$ goes from $1,...,N$, and $D_{N}$ is defined as
@@
    Generate 1000 random bootstrap samples from the assumed $F(.;\theta)$ , using the estimated parameters from Step 2. For each bootstrap sample:
@@
      Estimate the bootstrap parameters as in Step 2
@@
      Evaluate $Z_{i}$ using the estimated bootstrap parameters
@@
      Rearrange $Z_{i}$ in ascending order
@@
    Compare the original test statistic from Step 5 to the 95th percentile critical value of the bootstrap distribution. If it exceeds, we reject the null hypothesis
Theoretically, this procedure can be used when you are testing for any specified process completely determined by mean $\mu$ and variance $\eta^{2}$, see [3] for more details. We have only successfully tested it when the driving processes are Gamma, Inverse Gaussian, and Brownian motion, but as said above, we recommend that you use Procedure 1 when testing for Brownian motion, as it is easier computationally.
Under the $H_{0}$, we expect the p-value for test to be around 0.05. In table 8 below, we use Procedure 1 to test if the driving process is Brownian motion or not with varying $N$, $M$, and $a$ values. We refer to Section 2.2.1 of the supplemental material for the corresponding R code.
@@
In tables 9 and 10 below, we use Procedure 2 and test whether the driving process is some other specific process, i.e., Gamma in table 9 and Inverse Gaussian in table 10, under the $H_{0}$. We refer to Section 2.2.2 of the supplemental material for the corresponding R code.
@@
When the driving process is Gamma, we observe p-values below or very near 0.05 for each combination of $N$, $M$, and $a$ values.
@@
When the driving process is Inverse Gaussian, we observe p-values below or very near 0.05 for each combination of $N$, $M$, and $a$ values.
@@ 3.5.1 Power of the test
When the process is not driven by Brownian motion, i.e., Gamma, Inverse Gaussian, and mixed combinations of these, we expect high p-values for the test when testing for Brownian motion. In tables 11-13 below, the results for the power can be observed when using Procedure 1.
@@
When the driving process is Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
@@
When the driving process is Inverse Gaussian, we observe high p-values for each combination of $N$, $M$, and $a$ values.
@@
When the driving process is mixed Inverse Gaussian and Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
In order to illustrate the test’s effectiveness using Procedure 2, we observe the p-values for the test when testing for Brownian motion in tables 14-16 below, when using Procedure 2.
@@
When the driving process is Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
@@
When the driving process is Inverse Gaussian, we observe high p-values for each combination of $N$, $M$, and $a$ values.
@@
When the driving process is mixed Inverse Gaussian and Gamma, we observe high p-values for each combination of $N$, $M$, and $a$ values.
As a general comment, we observe a very good power of the test for all cases.
Although we include the above tables for only one case of the power, namely when the driving process is Brownian motion using Procedure 1 and Procedure 2, we have tested many combinations of these processes using Procedure 2. For example, when the driving process is Gamma and we test for Inverse Gaussian, and vise versa, the p-values are all similar to or greater than the rejection rates in the above tables.
We observe results that are consistent with our expectations both under the $H_{0}$, and for the power of the test. For different combinations of $N$, $M$ and $a$ values, we observe p-values around 0.05 under the $H_{0}$, and for the power of the test, we observe high p-values when testing for many different processes. The R code for this can be seen in Section 2.2 of the supplemental material.
4 Example with financial data

### v2

@@ Abstract
11footnotetext: Department of Mathematics and Statistics, Rhodes College, 2000 North Pkwy, Memphis, TN 38112, USA, Emails: abdelrazeqi@rhodes.edu
@@ 1 Introduction
Between 2014 and 2018, Abdelrazeq et al. developed theories and results indicating how to develop a model verification for Lévy-driven CAR(1) process [[1], [2], [3]]. Using these results, we develop a step-by-step methodological approach for how these findings can be used to verify the process and test its driving process distribution. We then show how to test the hypothesis that the CAR(1) process belongs to a specified class of Lévy processes. The performance of these test statistics under the null hypothesis is illustrated via simulation studies. To demonstrate the power of the test, we computed the power under one alternative case. We then illustrate how to apply the methodology step-by-step to many economic and financial data examples.
@@ 3 Methodology
In this section, we outline the step-by-step methodology for testing various background processes. The approach and results are clearly presented to enable users to apply these techniques to their own models and data. For further explanation of the methodology, along with the corresponding R code, refer to Sections 1 and 2 of the supplemental material.
@@ 3.1 Step 1: Choosing the right estimator for $a$
We are concerned with the consistency of the test statistic $W_{\Delta_{1}\widehat{L}^{(M)}}(1)$, defined when the parameter $a$ is replaced by an estimator..Therefore, one must first select the most accurate estimator for $a$.
@@
We need to consider an estimator for $a$ for the general second order driving process $L$ (i.e. $\mu\neq 0$). The following estimator of $a$:
$\widehat{{a}}_{N}^{(M)}=\frac{\sum_{n=1}^{NM}\left(Y_{\frac{n-1}{M}}-Y_{\frac{n}{M}}\right)\left(Y_{\frac{n-1}{M}}-\overline{Y}\right)}{\frac{1}{M}\sum_{n=1}^{NM}\left(Y_{\frac{n-1}{M}}-\overline{Y}\right)^{2}}~{}~{}\text{where}~{}~{}\overline{Y}=\frac{1}{NM}\sum_{n=1}^{NM}Y_{\frac{n}{M}},$ (3.1)
is based on the least squares approach, as defined and studied in [1]; hence, we will refer to it as the least squares-based estimator (LSB). This estimator is recommended when the driving process is an unspecified Lévy process or when Brownian motion is expected to be involved in the driving process. Additionally, it should be used to test whether the driving process is Brownian motion or not. The explanation and R code for this estimator can be found in Section 1.4.1 of the supplemental material.
@@
This estimator is based on the highly efficient Davis-McCormick estimator, as described in [7]; hence, we refer to it as the Davis-McCormick-based estimator (DMB estimator). It is highly accurate compared to the LSB estimator; however, the DMB estimator does not take negative values for $Y$. Therefore, if $Y$ is negative at any point, the LSB estimator should be used instead. Additionally, the DMB estimator should not be used when Brownian motion is expected to be involved in the driving process. We have tested and recommend this estimator for Gamma, Inverse Gaussian, Beta, and any mixed combinations of such Lévy processes. The explanation and R code for this estimator can be found in Section 1.4.2 of the supplemental material.
Below, in Tables 1-3, we compare the accuracy of our two estimators in simulated cases where the driving process is not Brownian motion. For these comparisons, we fix $\sigma=1$, $\mu=1$, $\eta=1$, $N=100$, $M=100$, and $\lambda=1$ in the relevant cases.
@@
For the simulated cases with Gamma, Inverse Gaussian, or Mixed Inverse Gaussian and Gamma-driven processes, we observe that the DMB estimator is more accurate in estimation for all values of $a$ than the LSB estimator. Although the DMB estimator is more precise than the LSB estimator in most simulated cases, both estimators are relatively accurate for each background process. Moreover, the LSB estimator should be used when the driving process is Brownian motion or any unspecified process.
@@ 3.2 Step 2: Choosing N and M
Before calculating the recovered increments, one should choose $N$ and $M$ such that $N/M$ is small, as discussed in [2]. $N$ can represent a day, a month, or a year, with $M$ being the number of observations in each period accordingly. Additionally, we have observed that for small values of $a$ and $N$, the test may encounter issues with empirical level (i.e., the approximated p-value) being close to the nominal level. Therefore, we recommend that $N$ be greater than $50$.
@@ 3.3 Step 3: Recovering increments with estimated $a$
To proceed, we consider the recovered increments using an estimator for $a$:
@@
See Section 1.7 of our supplemental material for the R code, along with a detailed explanation of how to calculate the recovered increments with the estimated $a$.
@@ 3.4 Step 4: Calculating the test statistic
We calculate $W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)$ as defined in Equation (2.8). For the significance level $\alpha=0.05$, we reject the null hypothesis of uncorrelated increments when $|W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)|>1.96$. In this case, we reject the null hypothesis that the data is Lévy-driven for every background process. If $|W_{\widehat{{\Delta_{1}\widehat{L}^{(M)}}}}(1)|<1.96$, we fail to reject the null hypothesis, and it is possible that a Lévy-driven CAR(1) process is a good model. In this case, we recommend proceeding to Step 5 (see subsection 3.5). Refer to Section 1.8 of the supplemental material for instructions on how to calculate the test statistic in R.
@@
Under simulated Lévy-driven CAR(1) processes, we now examine the performance of our verification process for various background processes (i.e., Brownian motion, Gamma, Inverse Gaussian, Beta, and mixed Lévy processes). The tables for each background process display the approximated p-value, for different values of $N$, $M$, and estimated $a$. See Section 2.1 of the supplemental material for the performance of the various background processes, along with the corresponding R code.
Note that a good indication of a test performing well under the null hypothesis is that, for a significance level of 0.05 and 400 simulations (i.e., $R=400$ in the following tables), the empirical level (i.e. the approximated p-value) should be around $0.05\pm 0.02$, which holds true in most of our cases. It is acceptable to have a deviation of 1 in 20, due to random chance, as predicted by the binomial distribution.
@@
For the Brownian motion-driven process, we observe the approximated p-values very close to $\alpha=0.05$ for most combinations of $N$, $M$, and $a$ values. It appears that the ratio $N/M$ has no significant effect. However, for $a=0.3$, and $a=0.5$, we observe that the approximated p-values are not very close to the nominal level. In this table, we use LSB estimator to estimate $a$, as recommended for Brownian motion and refer to [2] for further information on the effect of the LSB estimator on the results.
@@
For the Gamma and Inverse Gaussian driven process, we observe the approximated p-values very close to $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it appears that the ratio $N/M$ has no significant effect. In this table, we use the DMB estimator to estimate $a$, as recommended for these two cases.
@@
For the mixed Inverse Gaussian and Gamma driven process, Similarly, as above, we observe the approximated p-values very close to $\alpha=0.05$ for every combination of $N$, $M$, and $a$ values, and it appears that the ratio $N/M$ has no significant effect. In this table, we use the DMB estimator to estimate $a$, as recommended for this case.
Our results emphasize the strength of the verification process, as under the null hypothesis, we fail to reject the null hypothesis of uncorrelated increments for many Lévy-driven background processes, with approximated p-values very close to the nominal level, across varying combinations of $N$, $M$, and $a$ values. See Section 2 of the supplemental material for testing the performance of various background processes.
@@ 3.5 Step 5: Testing for the driving process
If the null hypothesis that the driving process is Lévy is not rejected, then the Lévy-driven CAR(1) could be considered a good model. In this case, the next step is to test whether the driving process is Brownian motion. Another interesting test is to determine whether the driving process is some other specific process, such as Gamma, Inverse Gaussian, or any other unspecified process. We recommend the following two procedures, generally adapted from [3] and [10], and modified for our verification process, to achieve these objectives.
@@
To test whether the driving process is Brownian motion (BM), we follow these steps.
@@
    Recover the estimated increments, as shown in Equation (3.3), using the LSB estimator.
@@
    Generate one bootstrap sample from the calculated estimated increments obtained in Step 1.
@@
    Compute the sample mean and standard deviation of the bootstrap sample obtained in Step 2.
@@
    Conduct the Kolmogorov-Smirnov (KS) test for the normal distribution using the sample mean and standard deviation calculated in Step 3.
This procedure should only be used if the process being tested is Brownian motion, as explained in [6] and [3]..
@@
To test whether the driving process is something other than Brownian motion, such as Gamma, Inverse Gaussian, or any other unspecified process, as outlined in [10], we follow these steps:
@@
    Recover the estimated increments, as shown in Equation 3.3, using the appropriate estimator (LSB or DMB).
@@
    Assuming the recovered increments follow a particular distribution $F(.;\theta)$, estimate the parameters for this distribution from the recovered increments.
@@
    Calculate the test statistic $D_{N}$ for the $N$ recovered increments $Z_{i}$, where where $i$ ranges from $1,...,N$, and $D_{N}$ is defined as
@@
    Generate 1000 random bootstrap samples from the assumed distribution $F(.;\theta)$, using the estimated parameters from Step 2. For each bootstrap sample:
@@
      Estimate the bootstrap parameters as in Step 2.
@@
      Evaluate $Z_{i}$ using the estimated bootstrap parameters.
@@
      Rearrange $Z_{i}$ in ascending order.
@@
    Compare the original test statistic from Step 5 to the 95th percentile critical value of the bootstrap distribution. If the original test statistic exceeds this value, reject the null hypothesis.
Theoretically, this procedure can be used when testing for any specified process completely determined by the mean $\mu$ and variance $\eta^{2}$, as detailed in [3]. However, we have only successfully applied it when the driving processes are Gamma, Inverse Gaussian, and Brownian motion. As mentioned earlier, we recommend using Procedure 1 when testing for Brownian motion, as it is computationally easier.
Under the null hypothesis $H_{0}$, we expect the estimated p-value for the test to be around the nominal level ($\alpha=0.05$). In Table 8 below, we use Procedure 1 to test whether the driving process is Brownian motion, with varying values of $N$, $M$, and $a$. The corresponding R code can be found in Section 2.2.1 of the supplemental material.
@@
In Tables 9 and 10 below, we use Procedure 2 to test whether the driving process is a specific process, namely Gamma in Table 9 and Inverse Gaussian in Table 10, under the null hypothesis $H_{0}$. The corresponding R code can be found in Section 2.2.2 of the supplemental material.
@@
When the driving process is Gamma or Inverse Gaussian , we observe p-values below or very near 0.05 for each combination of $N$, $M$, and $a$ values.
@@ 3.5.1 Power of the test
When the process is not driven by Brownian motion, such as Gamma, Inverse Gaussian, or mixed combinations of these, we expect high rejection rate (say $\hat{\beta}_{\triangle_{1}\hat{L}^{(M)}}$ for the test when testing for Brownian motion. In Tables 11-13 below, the results for the test power are presented when using Procedure 1.
@@
When the driving process is Gamma, Inverse Gaussian, or a mixed Inverse Gaussian and Gamma process, we observe a high rejection rate ($\hat{\beta}_{\triangle_{1}\hat{L}^{(M)}}$) for each combination of $N$, $M$, and $a$ values.
To illustrate the effectiveness of the test using Procedure 2, we observe the rejection rate ($\hat{\beta}_{\triangle_{1}\hat{L}^{(M)}}$) for testing Brownian motion in Tables 14-16 below, where Procedure 2 is applied.
@@
When the driving process is Gamma, Inverse Gaussian, or a mixed Inverse Gaussian and Gamma process, we observe a high rejection rate ($\hat{\beta}_{\triangle_{1}\hat{L}^{(M)}}$) for each combination of $N$, $M$, and $a$ values.
As a general observation, we find that both Procedure 1 and Procedure 2 exhibit very good power in all cases, with Procedure 2 having an advantage over Procedure 1. However, Procedure 1 is computationally easier, particularly when testing for Brownian motion.
Although the tables above show the power results for only one case-when the driving process is Brownian motion and we use Procedure 1 and Procedure 2-we have tested many other combinations of processes using Procedure 2. For example, when the driving process is Gamma and we test for Inverse Gaussian, and vice versa, the rejection rates are similar to or greater than those observed in the tables above.
We observe results that align with our expectations both under the null hypothesis ($H_{0}$) and for the power of the test. For various combinations of $N$, $M$ and $a$ values, the estimated p-values are around nominal level ($\alpha=0.05$) under $H_{0}$. For the test power, we observe high rejection rate when testing for a range of different processes. The corresponding R code can be found in Section 2.2 of the supplemental material.
4 Application to S&P 500 Stocks Dynamic Spread